An AM signal in which the carrier is modulated 70% contains 1500 W at the carrier frequency. Calculate the power content at the carrier when modulation percentage drops to 50%.
Answer #1

signal and Sali Am signal Afeter modulation = Yam(t) = Ac[1+ Kam14)] colorfcct where mit) is message Ae cos(irfet) is un moclwhen modulation percentage drops to 50% That is M 0.5 Pt2 = Pc [it Pe [1+ ] Pqz = 1209.819 (1+ 0:52] = 1904.819 [1 + 0,25) =| Hence, power content at the carrier is PT 2 = 1355.421 W when modulation percentage drops to 50%.